Hello, I am trying to solve this question... A pound of cashews sells for $6.00 and a pound of peanuts sells for $3.00. If you purchase a 4 pound mixture at the cost of $4.25 per pound, how many pounds of peanuts will you have? How many piunds of cashews?

Answer by checkley79(3341)

LET X BE THE AMOUNT OF CASHEWS & (4-X) BE THE AMOUNT OF PEANUTS.
6X+3(4-X)=4.25*4
6X+12-3X=17
3X=17-12
3X=5
X=5/3
X=1 2/3 POUNDS OF CASHEWS ARE USED.
4-(1 2/3)=2 1/3 POUNDS OF PEANUTS WERE USED.
PROOF:
6(1 2/3)+3(2 1/3)=4.25*4
10+7=17
17=17